Assume a par value of $1,000. Caspian Sea plans to issue a 7.00 year, semi-annual pay bond that has a coupon rate of 8.02%. If the yield to maturity for the bond is 7.66%, what will the price of the bond be?
Answer #1

Price of a bond is mathematically represented as: 1 1-- (1 + i) PEC. - +- м. (1 + i) where P is price of bond, with periodi

M = $1000, n = 7 * 2 = 14 semi-annual periods, C = 8.02% * $1000/2 = $40.1 (semi-annually), i = 7.66%/2 = 3.83% (semi-annually)

1- (1+0.0383)14 P = 40.1 * - 0.0383 1000 (1 + 0.0383) 14

P = $428.37 + $590.85

P = $1,019.23